OpenStudy (anonymous):

Is this a one to one function? F(x) = 1-x^2

OpenStudy (anonymous):

dont you have to set f(x1) = f(x2)?

OpenStudy (anonymous):

Are there any values, a and b, such that F(a)=F(b)? If that is the case, it is not one-one.

OpenStudy (anonymous):

no. 2 different x (+x & -x) corresponding the same F(x)

myininaya (myininaya):

function is a parabola upside down U so it is not 1 to 1

OpenStudy (anonymous):

ohh ok. so functions with even exponents are never one to one?

OpenStudy (anonymous):

heck no

OpenStudy (anonymous):

In this case, you can you 1 and -1, but that is just because it is an even function, it is not universal.

OpenStudy (anonymous):

ok thanks guys

OpenStudy (anonymous):

in order for a function to be one to one it must a) be strictly increasing or deceasing b) pass the "horizontal line test" (so parabolas are out)

OpenStudy (anonymous):

c) satisfy if \[f(a)=f(b) \] then \[a=b\]

myininaya (myininaya):

semicricles are out too!

OpenStudy (anonymous):

hello! do you like my new title?

myininaya (myininaya):

the only one to one circle is the kind with the radius 0 lol

myininaya (myininaya):

its okay satellite ( i helped you get there) i can go back through all the posts and take back medal i want to see if it will demote u lol

myininaya (myininaya):

jk satellite

OpenStudy (anonymous):

you can call me "master" now! lol

myininaya (myininaya):

master satellite when will you be guru